OpenStudy (anonymous):

What is the slope of the line that passes the given points? (6,-1) and (-3,-1)

OpenStudy (anonymous):

m= 0

OpenStudy (anonymous):

0 i believe

OpenStudy (azureilai):

the slope is zero, since the y-value between your poits never change. The graph would be y=-1

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

What is an equation of the line in slope intercept form?

OpenStudy (anonymous):

give us the equatiion

OpenStudy (azureilai):

2x+2

OpenStudy (anonymous):

no equation?

OpenStudy (azureilai):

y=mx+b, m=2, and b=2, so the equation would be y = 2x+2. according to the graph

OpenStudy (anonymous):

Correct Thanks i need help with 2 more.

OpenStudy (anonymous):

Write the equation in slope intercept form. What are the slope and y intercept? -9x + 10y= -9

OpenStudy (azureilai):

So in this equation you want y only in one side. So, -9x+10y=-9 10y=9x-9 (add 9x over) y= (9/10)x -9 (divide the whole thing by ten, and y is isolated) slope is the degree(change in y/change in x) that the line rises or drops and y intercept is where the line crosses the y axis

OpenStudy (anonymous):

is one of these?

OpenStudy (azureilai):

the last one. Notice that in the equation I solved, it is y = (9/10)x-9/10 (forgot to add "/10" to the second 9 earlier but this one is the correct one) since it is -9/10, the y intercept would be negative. and since m=9/10, the slope is positive. The answer is the last one

OpenStudy (anonymous):

Cool thanks 1 more question for now.

OpenStudy (azureilai):

Ok. but do you understand the process of solving your questions so far?

OpenStudy (anonymous):

What is the graph of the equation? -5x + y = -3

OpenStudy (anonymous):

or?

OpenStudy (azureilai):

m=5, and the y intercept is = -3. So from the graphs, which one do you think is correct?

OpenStudy (anonymous):

Which of these 4 answers would it be?

OpenStudy (anonymous):

c ?

OpenStudy (azureilai):

why do think it is c?

OpenStudy (azureilai):

You have the right idea, but you did not rearrange the equation. Remember, whenever you are graphing you always want to isolate y to one side of the equation. In this case, the equation becomes y= 5x-3, the slope is 5, and the y intercept is 3. the correct answer is a.

OpenStudy (azureilai):

y intercept is -3 *(forgot to add negative)

OpenStudy (anonymous):

Ohh yea thanks bud ill fan you and give you a medal.

OpenStudy (azureilai):

No problem.

OpenStudy (anonymous):

We already did this 1 can you please remind me which 1 was it.

OpenStudy (azureilai):

0

OpenStudy (anonymous):

Thanks